### v1

# DEEP-ICL: Definition-Enriched Experts for Language Model In-Context Learning

Abstract
It has long been assumed that the sheer number of parameters in large language models (LLMs) drives in-context learning (ICL) capabilities, enabling remarkable performance improvements by leveraging task-specific demonstrations. Challenging this hypothesis, we introduce DEEP-ICL, a novel task Definition Enriched ExPert Ensembling methodology for ICL. DEEP-ICL explicitly extracts task definitions from given demonstrations and generates responses through learning task-specific examples. We argue that improvement from ICL does not directly rely on model size, but essentially stems from understanding task definitions and task-guided learning. Inspired by this, DEEP-ICL combines two 3B models with distinct roles (one for concluding task definitions and the other for learning task demonstrations) and achieves comparable performance to LLaMA2-13B. Furthermore, our framework outperforms conventional ICL by overcoming pretraining sequence length limitations, by supporting unlimited demonstrations. We contend that DEEP-ICL presents a novel alternative for achieving efficient few-shot learning, extending beyond the conventional ICL.
DEEP-ICL: Definition-Enriched Experts for Language Model
In-Context Learning
thanks: Equal contributionXingwei Qu2,311footnotemark: 1Yiming Liang4,5,10 Yucheng Wang1 Tianyu Zheng1,9 Tommy Yue7Lei Ma8,10Stephen W. Huang11Jiajun Zhang4,5Wenhu Chen6 thanks: Corresponding author.Chenghua Lin322footnotemark: 2Jie Fu1,211footnotemark: 122footnotemark: 2Ge Zhang1,2,6 1M-A-P 2HKUST  3 University of Manchester 4Institute of Automation, CAS 5University of Chinese Academy of Sciences 6University of Waterloo 7Ohio State University 8Peking University 9Beijing University of Posts and Telecommunications 10BAAI 11Harmony.ai ge.zhang@uwaterloo.ca, jiefu@ust.hk, chenghua.lin@manchester.ac.uk
1 Introduction
Figure 1: Pipeline of DEEP-ICL. The preparation stage focuses on generating a pool of task-based experts, each trained using task-specific data. ICL firstly generates task definitions from given demonstrations using a task definition generator, which can be a custom generator, GPT, or human. Secondly, we utilize the task definitions to retrieve the task-based expert pool and match the top three experts that are most similar. Next, the expert ensemble step initializes the fusion expert by performing a weighted average of the matched experts’ weights. Finally, the data format is reconstructed for training the fusion expert and performing inference on the last input field.
In recent years, large language models (LLMs) have showcased remarkable capabilities, particularly for in-context learning (ICL). This method allows LLMs to learn how to execute specific tasks by analyzing a few labeled examples provided in their input context, without any updates to their parameters. This breakthrough enables LLMs to serve as versatile, general-purpose models capable of tackling a wide range of tasks with minimal guidance, as highlighted by Brown et al. (2020).
However, ICL also has its disadvantages. Most open-sourced LLMs such as LLaMA Touvron et al. (2023), Falcon Almazrouei et al. (2023), and ChatGLM Du et al. (2022); Zeng et al. (2022) have fixed input lengths, limiting the number of demonstrations that can be provided. Even with sufficient demonstrations, the performance of ICL cannot be enhanced. This limitation is due to ICL’s selective nature, requiring the choice of representative demonstrations, as proven by Levy et al. (2023).
In this paper, we propose a novel ICL-style ensembling approach through the lens of few-shot learning, called task Definition Enriched ExPert-ICL (DEEP-ICL). Inspired by Pan et al. (2023), our approach incorporates two 3B models: one focuses on task definition determination, and the other specializes in learning from task-specific demonstrations. Since few-shot learning usually targets specific tasks, it is difficult to generalize to other unseen tasks. To overcome this problem, we create a dynamic expert pool filled with prior knowledge during the demonstration learning phase. In other words, the experts in the pool can be selected according to the task definition derived in the task definition generation phase. Furthermore, this dynamic pool utilizes LoRA (Hu et al., 2021) for selecting the top experts based on task requirements. Our experiments show that even with continual training on just five samples, our ensembling methods outperform both traditional ICL and non-ensembling approaches. This setup supports lifelong learning by surpassing the fixed limitations of LLMs, significantly improving performance.
We evaluate DEEP-ICL on 117 subtasks in the official Super Natrual Instruction (SuperNI) Wang et al. (2022b) test dataset split Wang et al. (2022b). Our experiments, comparing the performance to LLaMA2 models with 7B and 13B parameters Touvron et al. (2023), and Falcon 7B Almazrouei et al. (2023), show that our models significantly outperform these two 7B level models and exhibit performance that is marginally below the LLaMA2-13B model, indicating our approach is competitive within a narrow margin. This result demonstrates that our solution is effective in addressing the limitations of ICL and few-shot learning techniques.
In conclusion, our contributions are as follows:
  • We propose a novel framework that delineates the roles between two models: task definition and task processing. This distinction further confirms that the primary challenge of in-context learning (ICL) is the extraction of task definitions, rather than the processing of tasks itself.
  • Through extensive experimentation, we have validated that the synergistic efficacy of our dual 3B model approach is on par with the ICL performance of the LLaMA2-13B model.
  • Our framework’s expert pool is designed to dynamically augment its database with input demonstrations, thereby enhancing and enriching its capacity on a continual basis.
2 Related Work
2.1 ICL and Few-Shot Learning with LLMs
The emergence of LLMs has revolutionized the landscape of machine learning, particularly for ICL and few-shot learning (Brown et al., 2020). These methodologies have been instrumental in enabling rapid model adaptation to new tasks, even with limited examples. This section delves into the intricacies of ICL and few-shot learning within the context of LLMs.
2.1.1 ICL in LLMs
ICL has gained significant attention with the rise of LLMs. ICL allows models to adapt to tasks based on contexts, without explicit gradient-based fine-tuning. A notable exemplar of this approach is GPT-3, which has demonstrated the capability to process tasks based on provided context Brown et al. (2020). This paradigm shift challenges traditional methods that rely on task-specific fine-tuning.
Further insights into ICL suggest that during their pre-training phase, LLMs internalize extensive array of “concepts”, positioning them within the domain of implicit Bayesian reasoning proposed by Xie et al. (2021). This perspective is further enriched by studies that draw parallels between LLMs and topic models, proposing that LLMs infer latent conceptual variables during their inference phase Wang et al. (2023).
A more granular understanding of ICL is provided by Pan et al. (2023) with characterizing two ways through task recognition (TR) and task learning (TL). While TR pertains to discerning tasks from demonstrations, TL is more oriented towards assimilating new input-label associations. This distinction underscores the intricate relationship between TR and TL, emphasizing the enhanced efficacy of TL when combined with model scaling and enriched demonstrations.
2.1.2 Few-shot Learning in LLMs
Few-shot learning aims to enable models to learn new tasks using limited annotated examples. LLMs, especially those pre-trained on a large scale like BERT Devlin et al. (2019), have showcased a remarkable ability to efficiently learn new tasks through fine-tuning Zhang et al. (2021). A notable approach in this domain is pattern-exploiting training, which reformulates text classification tasks into cloze questions or “prompts” that resemble masked language modeling Schick and Schütze (2020a, b). This approach has been further refined by techniques that automatically generate prompts and incorporate demonstrations into the input Gao et al. (2020), and by methods that densify the supervision signal with label conditioning Tam et al. (2021). However, because the downstream tasks are often learned in isolation, it is still difficult to generalize to completely unseen tasks. Ye et al. (2021) improves few-shot learning capabilities on unseen tasks by harnessing cross-task generalization ability from diverse known tasks.
2.2 Dynamic Adapter Composition Techniques
The ascent of Large Language Models (LLMs) has brought the necessity for agile and efficient adaptation techniques to the forefront, especially in light of the computational and memory demands inherent to these models. Two prominent strategies, Adapters and Mixture of Experts (MoE), have emerged as leading solutions in this space. While both champion modularity and adaptability, they offer distinct methodologies and advantages, which we explore in this section.
2.2.1 Adapters
Introduced by Houlsby et al. (2019), adapters have revolutionized neural architectures. These modular units merge well with pre-trained models, facilitating adaptation to novel tasks without tampering with the original parameters. This paradigm shifts from conventional fine-tuning, which extensively modifies the entire model, to adapters, which streamline the adaptation process by curtailing computational overhead. Furthering the adapter concept, low-rank adaptation (LoRA) Hu et al. (2021) freezes the base model parameters and trains a supplementary lightweight module instead.
Recent advancements in dynamic adapter composition include module decomposition and recomposition based on functionality Kingetsu et al. (2021), behavior modulation using task vectors Ilharco et al. (2022), and crafting modules through specific arithmetic operations Zhang et al. (2023). LoraHub Huang et al. (2023) stands out by leveraging the modularity of LoRA modules, facilitating the assembly of diverse modules.
2.3 Task Definition Extraction
For LLMs, task definition extraction is paramount, ensuring they generalize effectively across a wide array of tasks. a primary mechanism for task definition is through instructions. By providing models with lucid and precise instructions, they can better interpret and execute diverse tasks. This approach emerges as a versatile alternative to the traditional supervised learning paradigm. The efficacy of finetuning LLMs using instruction-output pairs has been demonstrated, marking significant progress in this domain Mishra et al. (2022); Yin et al. (2022); Sanh et al. (2022).
While the importance of high-quality instructions is undeniable, sourcing them often hinges on human expertise. This presents challenges in terms of scalability and uniformity. Some instruction sets, like OpenAssistant Köpf et al. (2023), are publicly available, but many remain proprietary. Addressing the intricacies of instruction generation, several innovative methodologies have been proposed. For instance, GPT-3 has been leveraged to derive instructions from initial seeds Honovich et al. (2022) autonomously. In a similar vein, the Self-instruct method Wang et al. (2022a) enables LLMs to generate both instructions and their associated outputs autonomously. Further enriching this domain, a novel technique has been introduced, which iteratively refines instruction prompts by integrating a seed model with a web corpus Li et al. (2023).
Our methodology, DEEP-ICL, aligns with the prevailing emphasis on instructions for task definition in LLMs. Unlike traditional methods focusing on autonomous instruction generation or refinement, DEEP-ICL integrates task definition extraction with expert ensembling. This approach ensures optimal model performance by accurately interpreting task definitions and distinguishing DEEP-ICL as a more efficient and adaptive solution in the LLM landscape.
3 Methods
DEEP-ICL represents a model ensembling method for generalizable and continual few-shot learning. This section delves into the details about how DEEP-ICL performs few-shot learning.
The first stage, Expert Pool Construction, focuses on generating task-based experts. These experts serve as knowledge candidates for our model. In the second stage, Task Definition Extractor, we introduce how to train a task definition extractor to conclude task definitions from given demonstrations. In the third stage, Task Definition Guided Retriever, the process of retrieving task experts based on input criteria is detailed. The fourth stage, Expert Ensembling, concentrates on ensembling tasks that correlate with the provided inputs. Finally, in the fifth stage, labeled Continual Few-shot Learning we explore the adaptability of our paradigm to support continual and lifelong learning, ensuring its flexibility and generalizability.
3.1 Expert Pool Construction
Within this section, our objective is to create an expert pool as our prior knowledge. Each expert specializes in a specific task and is trained with LoRA Hu et al. (2021). LoRA highlights the capability of pre-trained language models to exhibit a reduced dimension while maintaining their efficacy in training. Given $W_{0}$ from $\mathbb{R}^{d\times k}$, modification occurs by expressing the subsequent update as a low-rank decomposition, formulated as $W_{0}+\Delta W=W_{0}+BA$. Here, $B$ belongs to $\mathbb{R}^{d\times r}$, $A$ is from $\mathbb{R}^{r\times k}$, with the rank $r$ being substantially lesser than the lesser of $d$ and $k$. Throughout the training phase, $W_{0}$ does not update by gradient. However, $A$ and $B$ have parameters open for adjustments. Given $h=W_{0}x$, the adapted forward pass is:
$h=W_{0}x+\Delta Wx=W_{0}x+BAx$ (1)
For each task within SuperNI, we train a distinct expert LoRA adaptor. These adaptors can be integrated into our existing models. Consequently, for each task, we not only train a dedicated expert but also establish a tuple encompassing the instruction for each task and its corresponding adaptor. Given the context, we denote our expert pool by $\mathbf{E}$. Within $\mathbf{E}$, we encompass $N$ task units, each corresponding to a specific task from SuperNI. Each unit consists of a task definition, denoted as $\textbf{T}_{i}$, and its associated trained adaptor, represented as $\textbf{A}_{i}$.
3.2 Task Definition Extractor
In this section, we train a specialized task definition extractor. Given the in-context learning (ICL) format $\{\textbf{I}_{0},\textbf{L}_{0},\ldots,\textbf{I}_{k},\textbf{L}_{k},\textbf{I}_{\text{final}}\}$, the goal for language models (LLMs) is to produce the final label $\textbf{L}_{\text{final}}$. Here, $\textbf{I}_{i}$ and $\textbf{L}_{i}$ represent the given demonstration pairs input and label for $i=0,\ldots,k$, and $\textbf{I}_{\text{final}}$ is the final question for which we want the LLMs to respond with $\textbf{L}_{\text{final}}$. The parameter $k$ denotes the number of demonstrations. Our objective is to identify the format and task attribute similarities in the given demonstrations, ranging from $\textbf{I}_{0},\textbf{L}_{0}$ to $\textbf{I}_{k},\textbf{L}_{k}$, and to extract appropriate task definitions $\textbf{T}_{\text{concluded}}$ for them. Directly applying task definitions and demonstrations in ICL format into model training may result in issues with generalizing to out-of-distribution tasks. So, we align with the Stanford Alpaca framework Taori et al. (2023), add extra data in accordance with the Alpaca-GPT-4 instruction following data in Peng et al. (2023). This approach enables our extractor to summarize suitable task definitions for provided demonstrations with proper instructions.
3.3 Task Definition Guided Retriever
In this section, we elaborate on how to retrieve the relevant expert from the expert pool based on a given instruction. After acquiring an instruction, we input it into SentenceBERT(Reimers and Gurevych, 2019), resulting in an embedding $E_{\text{ref}}$. Similarly, all the task definitions within the pool are also processed through Sentence-BERT to generate their respective embeddings $E_{i}$. We then utilize the cosine similarity to compute the distances between $E_{\text{ref}}$ and each $E_{i}$. The formula for this distance metric is as follows: $\text{Distance}(E_{ref},E_{i})=\cos(E_{ref},E_{i})$ We subsequently select the top-$k$ most similar adaptors based on these computed distances and normalize each distance from $w_{0}$ to $w_{k}$ as our following model ensemble weights.
3.4 Expert Ensembling
Model Soup (Wortsman et al., 2022) is an ensemble technique that enhances performance by merging multiple models to serve as the initialization for the final model’s tuning. Inspired by this work, we consider our neural network defined as $\theta=\text{FineTune}(\theta_{0},w)$, where $\theta_{0}$ represents the parameters to be initialized with the adaptors from our expert pool $\mathbf{E}$, and $w$ denotes the ensemble weights discussed in Section 3.3. We employ a weighted average to initialize $\theta_{0}$ using the top-$k$ adaptors $\mathbf{A}_{1},\ldots,\mathbf{A}k$ as follows:
$\theta_{0}=\frac{1}{W}\sum_{i\in S}w_{i}\textbf{A}_{i},\quad\text{where}\quad W=\sum_{i\in S}w_{i}$ (2)
with $S\subseteq{1,\dots,k}$ being the set of indices corresponding to the top-$k$ adaptors adaptors as selected and discussed in Section 3.3.
3.5 Continual Few-shot Learning
After expert ensembling, we finetune our model using LoRA Hu et al. (2021) with the weight initialization $\theta_{0}$ and the given demonstrations $\{\textbf{I}_{0},\textbf{L}_{0},\ldots,\textbf{I}_{k},\textbf{L}_{k},\textbf{I}_{\text{final}}\}$. The final model parameters, $\theta_{\text{final}}$, are obtained as follows:
$\theta_{\text{final}}=\text{FineTune}(\theta_{0},w_{i}),$ (3)
where $\theta_{\text{final}}$ is used for inference on the final examples $\textbf{I}_{\text{final}}$ The benefits of our paradigm include the ability to handle demonstrations from any number of shots. Given a sufficient number of examples, we can further finetune the experts with these examples without limited token numbers like in ICL. Moreover, we can extend our expert pool $\mathbf{E}$, with the finetuned adaptor parameter $\mathbf{\theta}$ and the concluded task $\textbf{T}_{\text{concluded}}$ from the task definition extractor (as described in 3.2). To ensure diversity and uniqueness in expert pool $\mathbf{E}$, new task definitions $T_{\text{concluded}}$ are compared for similarity with existing entries $T$ in $\mathbf{E}$. If $T_{\text{concluded}}$ is sufficiently distinct (below a certain similarity threshold) from all entries in $\mathbf{E}$, it is added to the pool, thus updating $\mathbf{E}$ to include both diverse and unique definitions.
4 Experimental Setup
4.1 Evaluating the Performance of DEEP-ICL
Category All Coreference Resolution Data to Text Answerability Classification Question Rewriting Title Generation Word Analogy
K = 5 K = 500 K = 5 K = 500 K = 5 K = 500 K = 5 K = 500 K = 5 K = 500 K = 5 K = 500 K = 5 K = 500
T5-small 24.95 52.88 23.20 53.49 16.98 34.33 30.92 67.56 47.58 60.69 15.74 30.34 8.68 58.89
T5-base 37.17 64.75 32.13 65.82 26.05 40.42 55.10 77.46 55.13 63.52 22.82 36.53 12.04 99.94
T5-large 40.07 69.84 39.72 76.57 30.66 43.45 62.02 81.10 50.15 65.07 26.56 38.85 17.98 100.00
T5-3B 46.81 73.52 51.78 83.61 35.93 46.07 63.79 80.77 59.70 66.34 30.42 42.06 24.10 100.00
Category All Textual Entailment Cause Effect Classification Grammar Error Correction Overlap Extraction Keyword Tagging Dialogue Act Recognition
K = 5 K = 500 K = 5 K = 500 K = 5 K = 500 K = 5 K = 500 K = 5 K = 500 K = 5 K = 500 K = 5 K = 500
T5-small 24.95 52.88 22.95 59.10 35.63 53.84 62.94 78.4 17.62 59.07 25.55 50.39 26.18 62.83
T5-base 37.17 64.75 43.79 71.71 43.45 65.55 73.19 83.55 41.76 67.86 32.08 57.49 35.69 83.51
T5-large 40.07 69.84 42.52 80.17 49.08 74.02 69.42 83.39 36.44 70.07 36.03 61.19 39.33 89.62
T5-3B 46.81 73.52 50.28 88.04 53.63 79.04 82.99 85.96 43.72 73.30 43.39 64.70 47.58 90.45
Table 1: The performance of DEEP-ICL across different task categories. K represents the number of samples. Due to space constraints, we select scores from both the 5-sample and the 500-sample settings for result demonstration. For more details, please refer to Appendix Table 6.
We conduct experiments based on SuperNI, a dataset comprised of human-authored instructions, which encompasses a variety of tasks, to verify the generalization and performance enhancement of DEEP-ICL. We point out that only the parameters of the task definition extractor, task definition guided retriever, and ensembled expert are activated during the inference pipeline of DEEP-ICL. The estimated activated parameter size of DEEP-ICL is around 6B if we utilize our 3B paradigm. As a result, we select LLaMA2-7B, LLaMA2-13B, and Falcon-7B as our baselines, which the most widely utilized open-source models, demonstrating that DEEP-ICL achieves comparable few-shot performance to LLMs twice its size.
In this section, we detail the SuperNI-based experiment setting. Firstly, we construct an expert pool. For this purpose, we employ the T5 model series (Raffel et al., 2020) as our expert base model. Each T5 expert is specifically trained with LoRA (Hu et al., 2021) for each task from SuperNI official training splits to constitute DEEP-ICL’s prior knowledge. Due to many tasks being smaller in size, we opt for a 9:1 split for training and validation. Had we gone for an 8:2 train-test split, our overall training set would have been insufficient. For each task, adaptors and instructions are extracted from SuperNI. During training, the batch size is set to 40, the learning rate at 3e-4, the input length is 1024, and the LoRA trained parameters are the "q_proj" and k_proj modules in the multi-head attention modules.
Subsequently, we train a Task Definition Extractor using the Galactica 3B model. Within this framework, we utilized Alpaca-GPT-4 Peng et al. (2023) and 20K samples from SuperNI Wang et al. (2022b) for training. From each task, 200 samples were chosen, amounting to 15K in total. This shortfall is due to some tasks having less than 200 samples. To meet the requirement of 20K samples, we sampled more from tasks with a larger sample size. The model’s training follows the guidelines set by Stanford Alpaca, with a batch size of 32 and full parameter training.
During the Continual Training phase, we conduct a final round of continual learning, concentrating on the specified Input and Output. For model training, we utilized LoRA (as described by Hu et al. (2021)), focusing on the "q_proj" and "k_proj" modules within the multi-head attention mechanisms. The rank for LoRA was set to 4, the learning rate was established at 3e-4, and the input length was fixed at 1024.
4.2 Ablation Experiment
We conduct two ablation experiments to investigate the roles of the task definition and expert ensemble components in the DEEP-ICL process. We reconstructed the input-output data of demonstrations and created two formats of labeled data to train LoRA in two different ways.
The setting of Deep-ICL without definitions and ensemble, as stated in Table 2, refers to directly using few-shot input-output pairs as labeled data to train randomly initialized experts. In this scenario, the DEEP-ICL stage lacks both task definitions derived from user demonstrations and expert ensembling and degenerates into few-shot learning. We conduct these ablation experiments on settings with 5 samples, 50 samples, 100 samples, and 500 samples, as well as using T5 small, T5 base, T5 large, and T5 3B models, to investigate the impact of task definition in Deep-ICL under different data scales and model sizes.
The ablation experiment setting of Deep-ICL without ensemble firstly retrieves task definition from user demonstrations using our own 3B extractor. Next, we add the task definition into the input field in the user demonstrations and construct input-output pairs as labeled data for training LoRA. Compared to standard DEEP-ICL, DEEP-ICL without ensembling lacks the initialization of LoRA weights through expert ensembling and trains LoRA from scratch. We conduct the same experiments as DEEP-ICL without definitions and ensembling to explore the influence of expert ensembling in DEEP-ICL under different data scales and model sizes.
4.3 Analysing Task Definition Quality
We propose two methods to analyze the quality of task definitions concluded from user demonstration. One is an automatic quality analysis approach, and the other is a human analysis approach.
DEEP-ICL retrieves from the expert pool using the generated task definition and matches the top 3 experts based on similarity. This similarity is computed by using the task definition generated from user input and the task definitions stored in the expert pool. Automatic quality analysis approach counts the number of overlapping matched experts as a measure of the commonality of task definitions from different sources, such as task definitions generated by GPT-3, GPT-3.5, GPT-4, our extractor, and those stored in the expert pool.
We conduct human analysis on the quality of generated task definitions, comparing human task definitions with those generated by the model, and categorize them into the following four types: same, relevant task with missing details, relevant task with misleading information, different. Each sample in the human analysis is annotated by one of the authors, instead of crowd-sourcing annotators, to guarantee the validity of the human analysis.
5 Results
Figure 2: Performance comparison of DEEP-ICL across four distinct expert base model sizes and five task definition extractors with 5 demonstration inputs. The expert base models include LLaMA2-7B and LLaMA2-13B, and Falcon7B. Baseline ICL results are denoted by dashed lines. Notably, our paradigm employing a 6B parameters demonstrates superior performance to the LLaMA2-7B ICL and exhibits comparable results to the LLaMA2-13B ICL.
5.1 Performance of DEEP-ICL
As shown in Figure 2, our work introduces a novel task definition extractor, outperforms the LLaMA2-7B ICL while slightly trailing behind the LLaMA2-13B ICL outcomes. Futhermore, GPT-4 surpasses human-annotated results, particularly in contexts involving smaller model sizes. This result can be explained by the preference of human annotations to incorporate excessive detail into task definitions. Compared to GPT-4 short instructions, too many excessive constraints would mislead the attention of models. Despite our task definition extractor operating on a 3B parameter framework, it outperforms the GPT-3 model, which is based on a 175B parameter setup. This success is amplified by the application of supervised fine-tuning strategies, allowing our task generator extraction methodology to achieve enhanced results compared to those obtained through traditional GPT-3 implementations.
Figure 3: The Performances trend alongside increases in the number of demonstrations.
According to Figure 3 and 6, the increasing number of input demonstrations can significantly improve the accuracy of our models in each following categories, especially in smaller-sized configurations. For larger models, we observe a substantial performance improvement when increasing from 5 to 100 demonstrations. However, beyond 100 demonstrations, the rate of improvement in our model’s performance begins to plateau, indicating a more steady increase.
5.2 Ablation Study
Size 5 samples 50 samples 100 samples 500 samples
without definitions and ensembling (-44.71)
small 7.91 (-17.04) 7.79 (-22.38) 6.93 (-33.73) 7.41 (-45.47)
base 9.91 (-27.26) 9.79 (-40.25) 8.98 (-47.88) 8.51 (-56.24)
large 6.92 (-33.15) 7.75 (-52.03) 8.16 (-57.02) 8.63 (-61.21)
3B 8.57 (-38.24) 8.67 (-57.90) 8.91 (-60.64) 8.57 (-64.95)
without ensembling (-4.61)
small 13.51 ( -11.44) 17.10 ( -13.07) 37.10 ( -3.56) 51.50 ( -1.38)
base 26.40 ( -10.77) 47.10 ( -2.94) 54.40 ( -2.46) 64.59 ( -0.16)
large 25.32 ( -14.75) 57.12 ( -2.66) 63.27 ( -1.91) 68.88 ( -0.96)
3B 41.16 ( -5.65) 65.58 ( -0.99) 68.76 ( -0.79) 73.33( -0.19)
DEEP-ICL (w/ Trained TDG)
small 24.95 30.17 40.66 52.88
base 37.17 50.04 56.86 64.75
large 40.07 59.78 65.18 69.84
3B 46.81 66.57 69.55 73.52
Table 2: Ablation results. We conduct ablation experiments on T5_small, T5_base, T5_large, and T5_3B models, using 5 samples, 50 samples, 100 samples, and 500 samples. Olive color values and cyan color values indicate ablations of DEEP-ICL without definitions and ensembling and DEEP-ICL without ensembling, respectively. TDG refers to task definition generator.
Compared to the final DEEP-ICL, DEEP-ICL without definitions and ensembling exhibits an average performance decrease of 44.71% across various data scales and model sizes. This underscores the significant role of task definitions in DEEP-ICL, particularly within demonstrations of the same task type. In addition, another observation is that DEEP-ICL without definitions and ensembling does not show significant performance improvements as the data scale and model size increases, indicating that without explicit (ours) or implicit (ICL) task definition extraction, few-shot learning does not work. From another perspective, it underscores the crucial role of task definitions for this paradigm, both in ICL and in our approach.
In the ablation experiment setting of DEEP-ICL without ensembling, there are four experimental setups ([T5 small, 5 samples], [T5 small, 50 samples], [T5 base, 5 samples], [T5 large, 5 samples]) that perform more than 10% worse than the final DEEP-ICL. This indicates that in scenarios with small models or limited data, experts ensemble in DEEP-ICL facilitates rapid bootstrapping of inference for related tasks.
As the number of samples increases and the model size grows, the performance gap between final DEEP-ICL and DEEP-ICL without ensembling gradually diminishes. This suggests that the initialization of DEEP-ICL by using expert ensembling does not negatively impact the effectiveness of model finetuning in the future.
5.3 Task Definition Quality Analysis
5.3.1 Automatic Quality Analysis
Generator GPT3 GPT-3.5 GPT-4 Human Ours
GPT-3 351 25 16 14 10
GPT-3.5 25 351 98 55 41
GPT-4 16 98 351 76 40
Human 14 55 76 351 30
Ours 10 41 40 30 351
Table 3: The number of overlapping matched experts. The bold value represents the maximum count of overlapping matched experts between task definitions from different sources, while the underlined values represent the number of overlapping matched experts between human task definitions and other task definitions.
From Table 3, we observe that the overlap between the experts retrieved by the task definitions generated by our 3B generator and those retrieved by human task definitions is 30. This value is higher than the corresponding value of 14 in the GPT-3 experimental setting but lower than the corresponding values of 55 and 76 in the GPT-3.5 and GPT-4 experimental settings, respectively. This may be due to the lack of instruction fine-tuning in GPT-3, resulting in poorer instruction-following capabilities and thus generating lower-quality task definitions. In addition, GPT-3.5 and GPT4 match the highest number of the similar experts, which may be due to the semantic similarity in the task definitions they generate. Figure 4 illustrates the statistics of overlapping similar experts matched by pairwise task definition in detail.
5.3.2 Human Analysis
Number Same Sim_1 Sim_2 Different
GPT-3 0 10 1 106
GPT-3.5 30 29 24 34
GPT-4 47 51 15 4
Ours 9 22 19 67
Table 4: The number of categories. sim_1 corresponds to the category of relevant task missing details, while sim_2 corresponds to the category of relevant task with misleading information.
Table 4 reveals that the number of definitions generated by our 3B model that are the same or similar to the human task definition exceeds those of GPT3. But there is still a distance between our trained task definition extractor and GPT-3.5 and GPT4. This finding also corresponds to the conclusions of the automated quality analysis. GPT-3.5 and GPT4 based task definitions guided retrievers lead to lightweight but consistently better performance compared to our trained task generator. Please refer to Fig 5 for more details on manual analysis.
6 Conclusion
In this paper, we present DEEP-ICL, a new approach that addresses the limitations of existing ICL frameworks related to fixed input lengths and limited learning capabilities. DEEP-ICL utilizes an expert pool to provide a strong knowledge base and improve learning from input demonstrations. We divide the ICL process into task definition extraction and task processing. Our findings show that accurately extracting task definitions is vital for ICL success. During task processing, we employ a novel initialization technique based on task definitions, leveraging insights from the expert pool, resulting in significantly enhanced performance and advancing the ICL process.
Limitations
While our paradigm can alleviate the challenges associated with lifelong learning for Large Language Models (LLMs), the expansion of our expert pool with the addition of varied task definitions is inevitable. Nonetheless, a potential overlap in expertise within this pool may arise due to similarities between some experts. Addressing the optimization of this expert pool to manage such redundancies will be an essential focus of future work. In this section, an explicit task definition is extracted. Yet, whether a more effective approach could be achieved with solely implicit embeddings remains a subject for further discussion.
Ethics Statement
This paper does not raise any ethical concerns. The data and additional resources employed in this study are open-source and widely utilized in existing works.
Appendix A Auto Quality Analysis Appendix
Figure 4 illustrates the statistics of overlapping similar tasks matched by pairwise task definition., as detailed in Sections 5.3.1. We recorded the overlap count of experts retrieved for pairwise task definition queries and depicted them in a bar chart.
Appendix B Examples in Human Analysis
Figure 5 illustrates specific examples from our human evaluation process, as detailed in Sections 5.3.2. We categorize the samples into the following four classes: same, relevant task with missing details, relevant task with misleading information and different.
Appendix C The Table of Model Performance
Table 5 shows the performance of DEEP-ICL under different experimental settings, such as task definition generator, LoRA base model, and number of ICL examples. The human setting is to directly use the task definitions in the dataset.
Appendix D Model Performance Across Different Categories.
Table 6 displays the performance of DEEP-ICL across different task categories in detail.
Task Definition Generator LORA base Sample-5 Sample-50 Sample-100 Sample-500
GPT3 small 25.70 31.46 41.91 52.36
base 35.13 48.03 54.50 63.91
large 38.82 56.99 62.97 68.99
3B 44.74 64.74 69.03 73.51
GPT-3.5 small 29.06 32.33 42.44 53.72
base 38.36 50.17 57.20 64.49
large 41.77 60.32 65.81 70.61
3B 47.95 66.78 70.22 74.24
GPT4 small 30.47 36.11 42.22 52.14
base 39.09 49.59 56.67 64.74
large 43.14 60.27 66.00 70.68
3B 47.59 66.53 69.84 73.79
Ours small 24.95 30.17 40.66 52.88
base 37.17 50.04 56.86 64.75
large 40.07 59.78 65.18 69.84
3B 46.81 66.57 69.55 73.52
Human small 8.20 8.31 25.10 52.51
base 39.76 48.93 55.05 63.26
large 40.41 59.11 63.68 69.02
3B 47.48 65.62 68.71 71.75
No Definition small 7.91 7.79 6.93 7.41
base 9.91 9.79 8.98 8.51
large 6.92 7.75 8.16 8.63
3B 8.57 8.67 8.91 8.57
No Ensemble small 13.51 17.10 37.10 51.50
base 26.40 47.10 54.40 64.59
large 25.32 57.12 63.27 68.88
3B 41.16 65.58 68.76 71.33
Table 5: Performance of DEEP-ICL under different experimental settings, such as Task Definition Generator, LoRA base model, and number of ICL examples. The Human setting is to directly use the task definitions in the dataset. The No Definition setting is to directly train adapter soup without using task definition. The best results in each section are Bold, the second-best results are underlined.
(a) GPT3 & GPT-3.5
(b) GPT3 & GPT4
(c) GPT3 & Human
(d) GPT3 & Ours
(e) GPT-3.5 & GPT4
(f) GPT-3.5 & Human
(g) GPT-3.5 & Ours
(h) GPT4 & Human
(i) GPT4 & Ours
(a) GPT3 & GPT-3.5
(b) GPT3 & GPT4
(c) GPT3 & Human
(d) GPT3 & Ours
(e) GPT-3.5 & GPT4
(f) GPT-3.5 & Human
(g) GPT-3.5 & Ours
(h) GPT4 & Human
(i) GPT4 & Ours
(j) Ours & Human
Figure 4: The statistics of overlapping similar tasks matched by pairwise task definition. The horizontal axis represents the overlap count of similar tasks, while the vertical axis represents the number of matched tasks.
Figure 5: Examples of human analysis. We manually analyze the semantic similarity of paired instructions and categorize the samples into the following four classes: same, relevant task with missing details, relevant task with misleading information and different.
Category All Coreference Resolution Data to Text Answerability Classification
Sam 5 Sam 50 Sam 100 Sam 500 Sam 5 Sam 50 Sam 100 Sam 500 Sam 5 Sam 50 Sam 100 Sam 500 Sam 5 Sam 50 Sam 100 Sam 500
small 24.95 30.17 40.66 52.88 23.20 30.22 42.78 53.49 16.98 20.23 28.16 34.33 30.92 38.53 58.33 67.56
base 37.17 50.04 56.86 64.75 32.13 52.54 60.67 65.82 26.05 34.39 36.61 40.42 55.10 66.95 68.98 77.46
large 40.07 59.78 65.18 69.84 39.72 65.24 69.92 76.57 30.66 38.96 40.81 43.45 62.02 73.83 76.38 81.10
3B 46.81 66.57 69.55 73.52 51.78 69.30 73.93 83.61 35.93 42.80 45.10 46.07 63.79 75.99 77.84 80.77
Category All Question Rewriting Title Generation Word Analogy
Sam 5 Sam 50 Sam 100 Sam 500 Sam 5 Sam 50 Sam 100 Sam 500 Sam 5 Sam 50 Sam 100 Sam 500 Sam 5 Sam 50 Sam 100 Sam 500
small 24.95 30.17 40.66 52.88 47.58 51.11 57.24 60.69 15.74 19.66 23.14 30.34 8.68 9.05 18.29 58.89
base 37.17 50.04 56.86 64.75 55.13 59.96 61.38 63.52 22.82 32.20 33.77 36.53 12.04 36.08 69.68 99.94
large 40.07 59.78 65.18 69.84 50.15 62.80 65.03 65.07 26.56 35.22 37.06 38.85 17.98 80.71 97.81 100.00
3B 46.81 66.57 69.55 73.52 59.70 65.09 65.66 66.34 30.42 38.33 40.33 42.06 24.10 90.54 98.20 100.00
Category All Textual Entailment Cause Effect Classification Grammar Error Correction
Sam 5 Sam 50 Sam 100 Sam 500 Sam 5 Sam 50 Sam 100 Sam 500 Sam 5 Sam 50 Sam 100 Sam 500 Sam 5 Sam 50 Sam 100 Sam 500
small 24.95 30.17 40.66 52.88 22.95 30.28 43.58 59.10 35.63 41.87 52.52 53.84 62.94 65.73 71.59 78.40
base 37.17 50.04 56.86 64.75 43.79 56.09 63.50 71.71 43.45 54.97 56.92 65.55 73.19 80.52 81.16 83.55
large 40.07 59.78 65.18 69.84 42.52 64.00 72.32 80.17 49.08 60.85 65.01 74.02 69.42 82.01 84.40 83.39
3B 46.81 66.57 69.55 73.52 50.28 79.75 82.54 88.04 53.63 71.44 74.67 79.04 82.99 86.21 86.58 85.96
Category All Overlap Extraction Keyword Tagging Dialogue Act Recognition
Sam 5 Sam 50 Sam 100 Sam 500 Sam 5 Sam 50 Sam 100 Sam 500 Sam 5 Sam 50 Sam 100 Sam 500 Sam 5 Sam 50 Sam 100 Sam 500
small 24.95 30.17 40.66 52.88 17.62 27.70 39.24 59.07 25.55 30.71 34.41 50.39 26.18 28.77 43.71 62.83
base 37.17 50.04 56.86 64.75 41.76 56.40 62.40 67.86 32.08 42.72 50.78 57.49 35.69 55.12 68.81 83.51
large 40.07 59.78 65.18 69.84 36.44 62.71 66.27 70.07 36.03 54.18 56.51 61.19 39.33 69.61 82.26 89.62
3B 46.81 66.57 69.55 73.52 43.72 64.74 66.25 73.30 43.39 58.63 60.18 64.70 47.58 78.84 84.54 90.45
Table 6: The performance of DEEP-ICL across different task categories.